Find the area that us inside r=4 and outside r=4-2sinθ

1
Expert's answer
2022-06-10T10:45:38-0400
"4=4-2\sin \theta"

"\theta_1=0, \theta_2=\pi"

"A=\dfrac{1}{2}\displaystyle\int_{0}^{\pi}\bigg((4)^2-(4-2\sin \theta)^2\bigg)d\theta"

"=\dfrac{1}{2}\displaystyle\int_{0}^{\pi}\bigg(16\sin \theta-4\sin ^2\theta\bigg)d\theta"

"=\dfrac{1}{2}\displaystyle\int_{0}^{\pi}\bigg(16\sin \theta-2+2\cos 2\theta\bigg)d\theta"

"=[-8\cos \theta-\theta+\dfrac{1}{2}\sin 2\theta]\begin{matrix}
 \pi\\
 0
\end{matrix}"

"=8-\pi+0-(-8-0+0)"

"=16-\pi ({units}^2)"